the sum of the speeds of two trains is726.6 miles per hour. the speed of the fist train is 9.4 mph faster than the second train. find the speed of each train
Answer by nerdybill(7384)

the sum of the speeds of two trains is726.6 miles per hour. the speed of the fist train is 9.4 mph faster than the second train. find the speed of each train
.
Let x = speed of slower train
then
x+9.4 = speed of faster train
.
From:"sum of the speeds of two trains is 726.6 miles per hour" we get our equation:
x + (x+9.4) = 726.6
2x+9.4 = 726.6
2x = 717.2
x = 358.6 mph (slower train)
.
faster train:
x+9.4 = 358.6+9.4 = 368 mph